Find the sum of all the possible three digit numbers formed by using the digits 1,2,3(repeatiton allowed).

Grade:12

1 Answers

Saurabh Koranglekar
5 years ago
numbers 1,2,3 would occur 9 times at each hundred, tens and units place

hence the sum of all the 27 numbers =
[(1+2+3) * 100 + (1+2+3) * 10 + (1+2+3) * 1 ] * 9

= 6 * 111 * 9 = 54 * 111 = 5994.

hope it helps otherwise revert